OpenStudy (anonymous):

7a=2-b 7a=5-b use the method of substitution to solve the system of linear equation. Please HELP!!!!!

OpenStudy (barrycarter):

Well, if both 2-b and 5-b are equal to 7a, then they're equal to each other, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

its asking me if there is a solution..if so what is the ordered pair

OpenStudy (anonymous):

2-b=5-b add a b to both sides 2=5 what does this mean?

OpenStudy (anonymous):

substitution?

OpenStudy (mathmate):

This is a substitution. Some people call it a comparison.

OpenStudy (anonymous):

there in this equation there is infinite number of solutions right?

OpenStudy (anonymous):

wrong 2=5 2 cannot equal 5 thus meaning?

OpenStudy (anonymous):

there is no solution

OpenStudy (anonymous):

right if the final solution was 2=2 this would mean there are infinitely many solutions a=2 means there's one solution and you would need to find the other coordinate 5=2 means there are no solutions simply due to the fact that it is an incorrect statement

OpenStudy (anonymous):

ty